Which of the followings is not used in forecasting based on the simple exponential smoothing method?
A) The most recent forecast for the past year
B) Precise actual demand for the past year
C) The value of the smoothing constant
D) Trend for the past year

Please explain.
Answer #1

Answer: (D) Trend for the past year

Explanation: Simple exponential smoothing, as the name suggests, is the simplest form of all the exponential smoothing forecasting methods. It is used when there is no clear trend or seasonal patterns in the given data set.

In simple exponential smoothing,

F(t+1) = A(t) + (1-) F(t)

where,

F(t+1) = Forecast for the next period

A(t) = Actual demand for the past year

F(t) = Forecast for the past year

= Smoothing constant